Sam has 6 rose bushes. He counted the flowers on each of them. There are 8, 2, 5, 4, 11 and 9. Find the Standard Deviation. Is X a “Usual” number of flowers? X is your last two digits of your CUNYFirst ID Number. (00 = 0, 01 = 1, …) (20 points) my id is 25


1
Expert's answer
2022-03-25T05:24:12-0400

"EX=\frac{8+2+5+4+11+9}{6}=6.5\\EX^2=\frac{8^2+2^2+5^2+4^2+11^2+9^2}{6}=51.8333\\DX=EX^2-\left( EX \right) ^2=51.8333-6.5^2=9.5833\\\sigma X=\sqrt{DX}=\sqrt{9.5833}=3.09569\\25 is\,\,not\,\,a\,\,usual\,\,number, it\,\,is\,\,not\,\,in\,\,the\,\,range\,\,of\,\,X"